# SmileyLlama: Modifying Large Language Models for Directed Chemical Space Exploration

### v3

@@ 3.4 SmileyLlama Outside of Chemical Language Modeling
While SFT and DPO clearly alters Llama in the creation of SmileyLlama, one might wonder whether SmileyLlama still retains the full LLM knowledge of the Llama-3.1-8B model from which it derives. We find that SmileyLlama can still converse in English if it is prompted to do so, and some sample conversations are included in Section 3 in the Supplementary Information. However, we notice that its responses are often shorter than those of Llama-3.1-8B and can contain SMILES string non-sequiturs, which seem to be more likely when dealing with prompts relating to chemistry. As a more quantitative measure of its residual abilities, we evaluate performance of SmileyLlama on the Language Model Evaluation Harness containing the MMLU GPQA, Math-Hard, and MMLU-Pro benchmarks74, 75, 76, 77, 78. Supplementary Table S3 and Figure S4 show that SmileyLlama generally performs worse on moral scenarios, and amusingly SmileyLlama also performs worse on subjects dealing with chemistry than Llama. This in part is likely due to the tendency for SmileyLlama to complete prompts relating to chemistry with a SMILES string as noted above. In addition, accuracy errors in the MMLU tests have also been noted recently79, and thus SmileyLlama’s degraded performance in chemistry may also be in part an artifact of poor testing benchmarks. Overall this result is somewhat encouraging, since it implies the possibility that LLM-derived CLMs can inherit and take advantage of the natural language processing ability of their foundation model. SmileyLlama already does this— we can steer the properties of the molecules it generates and the chemical space it explores using natural language prompts while still retaining some ability to process non-chemical natural language. However, more work will be required to view SmileyLlama as an additional capability of an LLM, as opposed to its current and primary benefit as a CLM.
@@
We have investigated the adaptation of LLMs to molecular chemistry problems that have been largely in the domain of CLMs. We find that the key advantage that LLMs have when performing as CLMs comes from prompting: information about the region of interest in chemical space can be specified in the prompt, letting the model generate chemical structures in this region. Furthermore, we showed that training an LLM on a variety of prompts allows for good performance in multiple CLM tasks, and found evidence to suggest that the LLM models can correctly generalize to scenarios with new prompts that leverage the abilities developed when trained on similar prompts. Concretely, we showed that Meta-Llama-3.1-8B-Instruct can be converted into a CLM on par with or exceeding the performance of modern CLMs7, 8, 19, 20, 21, rather than only demonstrating that LLMs can modify currently existing SMILES strings to have different properties.
@@
Our study clarifies a few crucial points for CLMs derived from LLMs going forward. First, it is not necessary to pretrain a specialized model on chemistry-specific text to generate molecules from a text description; a much less resource-intensive SFT training run on prompt-following on a dataset of a few million molecules with a commodity LLM is enough to do that. Second, DPO provides another resource-efficient way of optimizing the model to produce molecules that score well on a targeted objective without needing in-context examples, instead relying on the generative nature of the model itself for good and bad examples. Using these optimization methods, we showed that data specific to a target protein can generate increasingly better inhibitor molecules as determined by docking scores. A surprising corollary to this is the finding that SmileyLlama can combine its knowledge gained during single-objective optimization to perform well at a task specifying multiple objectives, elicited by combining the prompts. SmileyLlama was able to generate molecules which satisfy Lipinski’s rule of 5 and a high docking score on SARS2-Mpro even though SmileyLlama was not trained on the generation of molecules with both prompts, but instead is found to generate molecules with both attributes at inference.
@@
The prompting and optimization framework for modifying LLMs to explore chemical space broadly or to narrow the search to specific regions shown here could also be leveraged for molecular design outside of drug discovery, such as the use of SMILES for elaborating on transition metal complexes80. One could also imagine that casting a chemical problem as a linguistic construct could also allow for new applications, such as our recent work on chemical synthesis41. As with many of the fields touched by LLMs this decade, the newly opened frontier of possibility in chemistry is as vast as it is exciting.
@@ Acknowledgments
Code used for this study can be found at https://anonymous.4open.science/r/SmileyLlama-gh-D038/
@@
Supplementary Information
SmileyLlama: Modifying Large Language Models for Directed Chemical Space Exploration
Joseph M. Cavanagh1, Kunyang Sun1, Andrew Gritsevskiy2, Dorian Bagni1, Thomas D. Bannister3,
Teresa Head-Gordon1,4,5†
1Kenneth S. Pitzer Theory Center and Department of Chemistry,
University of California, Berkeley, CA, 94720 USA
3Department of Molecular Medicine, The Herbert Wertheim UF Scripps Institute for Biomedical Innovation and Technology, 130 Scripps Way, Jupiter, FL, 33458
University of California, Berkeley, CA, 94720 USA
$\dagger$ thg@berkeley.edu
S1 Details of properties to fine-tune
Overview of selected properties for fine-tuning. When fine-tuning Llama to generate drug-like molecules, we carefully assess various design choices and proceed with the following properties, emphasizing those that medicinal chemists would consider when proposing de novo drug molecules. We categorized and summarized all 12 properties into 4 subgroups as follows.
  • Physiochemical properties. Absorption, distribution, metabolism, and excretion (ADME) are the crucial criteria to quantify the localization and concentration of drug molecules within the body after administration. As a result, we build on the list of properties proposed in the classical Lipinski’s rule of 5 62 with some modern additionstpsa, WEI20201839 to generate drug-like molecules that could demonstrate decent ADME.
    • Number of Hydrogen Bond Donors (#HBD)
    • Number of Hydrogen Bond Acceptors (#HBA)
    • Molecular Weight (MW)
    • Log of Partition Coefficient (LogP)
    • Topological Polar Surface Area (TPSA)
    • Fraction of $sp^{3}$ hybridized carbon atoms ($Fsp^{3}$)
  • Structure flexibility features. Binding sites within a targeted biomolecule (most often a protein) display by nature complex 3D geometry, with key potential sites of drug-target interactions (amino acid side chains, as an example) somewhat fixed in space. The protein, however, has a dynamic structure and even the binding pocket undergoes changes in shape. Drug-like molecules need to be sufficiently rigid to make efficient interactions with the target protein, including in most cases a high degree of selectivity over making corresponding interactions with related proteins). Perhaps less intuitive is that drug-like molecules must be flexible enough to maintain those interactions as the protein adapts its conformation. There is a "Goldilocks principle" at play, where too rigid or too flexible are each undesired extremes. Here, we chose the following two properties to account for the flexibility aspect.
    • Number of Rotatable Bonds (#rot)
    • Whether the molecule contains a macrocycle (defined as an 8-membered ring or larger)
  • Pattern-based features. In practical drug discovery, there are always some key patterns and/or scaffolds that medicinal chemists would like to hold onto or get rid of. For instance, in the lead optimization phase, retaining the key moiety and desired chemical formula are rather essential. Meanwhile, avoiding chemically unstable groups, PAINS molecules baell_chemistry_2014, and molecules that would cause structure alerts could increase the chance of success in development. Therefore, we have the following three properties for fine-tuning.
    • Avoidance of undesirable chemical patterns
    • Retention of specified substructure (between 50 and 250 Da in molecular weight)
    • Chemical formula
  • Covalent warhead feature. Drugs can be broadly categorized into noncovalent and covalent drugs, depending on whether the drug reacts with its target. For example, a the electrophilic group of a covalent inhibitor might form a bond with a nucleophilic amino acid side chain of its target protein. The reactive functional group of a covalent inhibitor is called a warhead. While most drugs are non-covalent, either can be desired. To give the model the ability to generate covalent binders, we also curated common covalent warhead-related SMARTS patterns from the Enamine fragment library 61 to indicate whether our generated molecules have the capacity to covalently bind to the target or not.
    • Whether the molecule contains common covalent warhead-related SMARTS patterns, and which of these patterns appear in the molecule
Prompting options used in fine-tuning. To incorporate the properties mentioned above into the training, we used several ways of prompting to satisfy the requirement from target uses.
For numerical properties, including all physiochemical properties and #rot, we prompted Llama by providing a specific range that the training molecules falls into for that specific category. All the cutoff values used for ranges are either commonly used standards in drug discovery or derived from the training distribution. Besides the range guidance, we added the prompt that tell Llama exactly how many #HBDs and #HBAs are contained in the training data that gives its ability to do more nuanced generation. If a property falls into multiple valid ranges—for instance, 4 H-bond donors is less than or equal to all of 4, 5 and 7—we select one of the ranges at random to place in the prompt (if the properly is picked to be placed into the prompt). It is important that the list of ranges for a property span all possible molecules, otherwise a prompt which omits information biases the model toward producing molecules with values of that property outside of the list of ranges. If we never included information in the prompt about when a molecule had over 7 H-bond donors, but did sometimes include information about the number of H-bond donors when they number under or equal to 7, then omitting the information would give the model a hint that the number of H-bond donors is more likely to be over 7. Doing this during training would bias results during inference. This is the same reason we specify undesirable properties such as the presence of bad SMARTS strings in the prompt sometimes. If the random number generator decided that a prompt should contain a substructure but the SMILES in question did not have any BRICS substructures, we added "no BRICS substructure" to the list of properties in lieu of a substructure.
For other categorical properties, we used a combination of RDKit modules, SMILES strings, and SMILES arbitrary target specification (SMARTS) strings to recognize if certain properties or chemical patterns are present in the training input. Unlike the objective of containing the scaffold exactly, chemical pattern avoidance and covalent warhead recognition required matching of more general substructures and/or certain functional groups. Here, we used SMARTS strings as our representation because of its ability of matching chemical patterns. More details about the specific SMARTS patterns used are shown later in this section.
Below is a detailed list of possible components of that could appear in a training prompt.
  • $N$ H-bond donors, $N=\leq 3,\leq 4,\leq 5,\leq 7,>7$
  • $N$ H-bond acceptors, $N=\leq 3,\leq 4,\leq 5,\leq 10,\leq 15,>15$
  • $N$ Molecular weight, $N=\leq 300,\leq 400,\leq 500,\leq 600,>600$
  • $N$ LogP, $N=\leq 3,\leq 4,\leq 5,\leq 6,>6$
  • $N$ Rotatable bonds, $N=\leq 7,\leq 10,>10$
  • $N$ Fraction $sp^{3}$, $N=<0.4,>0.4,>0.5,>0.6$
  • $N$ TPSA, $N=\leq 90,\leq 140,\leq 200,>200$
  • a macrocycle, no macrocycles
  • has bad SMARTS, lacks bad SMARTS
  • has covalent warheads, lacks covalent warheads
  • substructure of *a_smiles_string*
  • a chemical formula of *formula*
SMART patterns used to identify bad chemical groups. Li et al. pointed out a list of bad chemical patterns that exists in ChEMBL database, which will negatively affect compound generation21. In this work, we used the same list of SMARTS patterns as their work to avoid bad patterns, including cyclopentadiene, cyclopentadiene ylidenes, aromaticity-breaking tautomers, antiaromatic system, unstable halogen-heteroatom bonds, unstable fused rings, allenic system, thiazyl linkages, and peroxide bonds. In Table S1, we also present the frequency of sampling undesirable chemical groups in ChEMBL and across different generative models.
  • [C^2]1=[C^2]-[C^2]=[C^2]~[C;!d4]~[C;!^2;d2]1
  • [C^2]1~[C^2]~[C^2]~[C^2]~[C;!^2;d2]~[N]1
  • [#6^2]1~[#6^2]~[#6^3;!d4]~[#6^2]2~[#6^2]~
    [#6^2]~[#6^2]~[#6^2](~[*])~[#6^2]~2~[#6^2]~1
  • [#6]1(=[*])[#6]=[#6][#6]=[#6]1
  • [#6]1=[#6][R{2-}]=[R{2-}]1
  • [#6^2]1~[#6^2]~[#6^2]~[#6^2]~[#6^1]~[#6^1]~1
  • [#7,#8,#16]-[#9,#17,#35,#53]
  • [r3,r4]@[r5,r6]
  • [*]=[#6,#7,#8]=[*]
  • [#7,#16]=[#16]
  • [#8]-[#8]
In addition to the patterns mentioned above, we use the following SMARTS patterns to enforce our generated pyrroles to be one of the following correct forms.
  • [N^2]1~[C,N;^2](=[*])~[C,N;^2]~[C,N;^2]~[C^3]1
  • [N^2]1~[C,N;^2]~[C,N;^2](=[*])~[C,N;^2]~[C;^3]1
  • [N^2]1~[C,N;^2]~[C,N;^2]~[C,N;^2](=[*])~[C;^3]1
  • [C,N;^2](=[*])1~[N;^2]~[C,N;^2]~[C,N;^2]~[C;^3]1
  • [C,N;^2]1~[N;^2]~[C,N;^2](=[*])~[C,N;^2]~[C;^3]1
  • [C,N;^2]1~[N;^2]~[C,N;^2]~[C,N;^2](=[*])~[C;^3]1
SMART patterns used to encode common covalent warhead-related functional groups. Common covalent warheads are extracted from the Enamine Covalent Screening and Covalent Fragment Library enamine_covalent_libraries. The list of SMARTS strings is shown below.
  • sulfonyl fluorides: [#16](=[#8])(=[#8])-[#9]
  • chloroacetamides: [#8]=[#6](-[#6]-[#17])-[#7]
  • cyanoacrylamides: [#7]-[#6](=[#8])-[#6](-[#6]#[#7])=[#6]
  • epoxides: [#6]1-[#6]-[#8]-1
  • aziridines: [#6]1-[#6]-[#7]-1
  • disulfides: [#16]-[#16]
  • aldehydes: [#6](=[#8])-[#1]
  • vinyl sulfones: [#6]=[#6]-[#16](=[#8])(=[#8])-[#7]
  • boronic acids/esters: [#6]-[#5](-[#8])-[#8]
  • acrylamides: [#6]=[#6]-[#6](=[#8])-[#7]
  • cyanamides: [#6]-[#7](-[#6]#[#7])-[#6]
  • chloroFluoroAcetamides: [#7]-[#6](=[#8])-[#6](-[#9])-[#17]
  • butynamides: [#6]#[#6]-[#6](=[#8])-[#7]-[#6]
  • chloropropionamides: [#7]-[#6](=[#8])-[#6](-[#6])-[#17]
  • fluorosulfates: [#8]=[#16](=[#8])(-[#9])-[#8]
  • beta lactams: [#7]1-[#6]-[#6]-[#6]-1=[#8]
In section 3.2, we investigate SmileyLlama’s performance on the following 387 tasks, grouped into the families for which the averages are shown in the table.
  • exactly k H-bond donors, from k=0 to k=5
  • exactly k H-bond acceptors, from k=0 to k=10
  • <= k H-bond donors, for k=3,4,5,7
  • <= k H-bond acceptors, for k=3,4,5,10,15
  • <= k Molecular weight, for k=300,400,500,600
  • <= k LogP, for k=3,4,5,6
  • <= 7, <=10, >10 Rotatable bonds
  • > 0.4, > 0.5, > 0.6, < 0.4 Fraction sp3
  • <=90, <=140, <=200 TPSA
  • a macrocycle
  • no macrocycles
  • has bad SMARTS (not shown in table but included for completeness)
  • lacks bad SMARTS
  • lacks covalent warheads
  • has covalent warheads (one for each of the 16 covalent warheads in the section above)
  • a substructure of (one of each of 320 Enamine fragments 61)
  • <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 Molecular weight, <= 5 LogP
  • <= 3 H-bond donors, <= 3 H-bond acceptors, <= 300 Molecular weight, <= 3 LogP
S2 Algorithms used for SFT and DPO
Algorithm 1 Pseudocode for generating the SFT dataset
  for all SMILES_STRING $\in$ ChEMBL do
     PROMPT $\leftarrow$ "Output a SMILES string for a drug like molecule with the following properties:"
     for all (PROPERTY_FUNCTION, PROPERTY_NAME) $\in$ shuffle(PROPERTIES) do
        VALUE $\leftarrow$ PROPERTY_FUNCTION(SMILES_STRING)
        if random_coin_flip() $=$ HEADS then
           PROMPT $\leftarrow$ PROMPT $+$ PROPERTY_NAME $+$ " of " $+$ VALUE $+$ ", "
        end if
     end for
     prompts.append(PROMPT)
     completions.append(SMILES_STRING)
  end for
Algorithm 2 Pseudocode for generating the DPO dataset for improving prompt obedience
  for all (SMILES_STRING, CONDITIONS) $\in$ RESPONSES do
     if MEETS_CONDITIONS(SMILES_STRING, CONDITIONS) $=$ True then
        winners.append(SMILES_STRING)
     else
        losers.append(SMILES_STRING)
     end if
  end for
  pairs $\leftarrow$ randomly_pair(winners, losers)
  assert: $|\text{pairs}|=\min\{|\text{winners}|,|\text{losers}|\}$
Algorithm 3 Pseudocode for generating the DPO dataset for Mpro optimization every iteration
  PROMPT $\leftarrow$ "Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO"
[…]
     SMILES_LIST $\leftarrow$ SmileyLlama.inference(PROMPT, num_responses)
     SMILES_LIST $\leftarrow$ SMILES_LIST.remove_redundancies()
  end while
  for SMILES $\in$ SMILES_LIST do
     iMiner.compute_score(SMILES)
  end for
  for i $=$ 1 to 8 do
     for SMILES $\in$ SMILES_LIST do
        while True do
           RANDOM_OTHER_SMILES $\leftarrow$ pick_random(SMILES_LIST)
           if iMiner.retrieve_score(SMILES) $>$ iMiner.retrieve_score(RANDOM_OTHER_SMILES) then
              DATASET.append((PROMPT, SMILES, RANDOM_OTHER_SMILES))
              break
           else if iMiner.retrieve_score(SMILES) $<$ iMiner.retrieve_score(RANDOM_OTHER_SMILES) then
              DATASET.append((PROMPT, RANDOM_OTHER_SMILES, SMILES))
              break
           end if
        end while
     end for
  end for
S3 SmileyLlama Outside of Chemical Language Modeling
While SFT clearly alters the model, one might wonder whether SmileyLlama retains LLM abilities unrelated to exploring chemical space. We find that SmileyLlama inherits some ability to process natural language, however, SmileyLlama will sometimes respond to chemistry-related queries with a SMILES string rather than a sensible, natural-language answer. Below are four sample responses from SmileyLlama to four prompts. Each was generated with greedy decoding, which is equivalent to sampling at a temperature of 0. All responses were generated using the Alpaca instruction format, with a system prompt/instruction of “You are a helpful assistant”. As shown below, SmileyLlama is capable of writing working python code for simple functions and answering some questions in natural language. This demonstrates that, while SmileyLlama was designed to be a CLM, it still inherits some of Llama’s natural language processing ability. However, this is not perfect, as SmileyLlama will sometimes respond to chemistry-related queries with a SMILES string, rather than the correct answer.
Input: Write me some python code which calculates the n’th Fibonacci number
Output:
def fibonacci(n):
if n <= 0:
return "Input should be a positive integer"
elif n == 1:
return 0
elif n == 2:
return 1
else:
a, b = 0, 1
for _ in range(2, n):
a, b = b, a + b
return b
Input: What is the molecular weight of hydrazine?
Output: N1N=C(C)C(C1)C
(Note: This is the SMILES representation of 3,4-dimethyl 2-pyrazoline, rather than hydrazine)
Input: What is a Diels-Alder reaction in organic chemistry?
Output: A Diels-Alder reaction is a [4+2] cycloaddition between a diene and a dienophile.
Input: Who was Christopher A. Lipinski in the field of chemistry? What was his "Rule"?.
Output: Christopher A. Lipinski was a chemist who developed the "Lipinski’s Rule of Five" which states that for a drug to be orally bioavailable, it should have the following properties:
1) C(C)Cn1c2c(cccc2)c(c1)C(Nc1ccccc1)=O
2) A molecular weight of 500 or less
3) No more than 5 H-bond donors
4) No more than 10 H-bond acceptors
Furthermore, we used the Language Model Evaluation Harness to test SmileyLlama’s performance on the MMLU GPQA, Math-Hard, and MMLU-Pro benchmarks 74, 75, 76, 77, 78. The results in Table S2 and Figure S3 show that SmileyLlama generally performs worse on most tasks compared to Llama, especially difficult tasks as assessed in the latter three benchmarks. For practical purposes, the SFT procedure for creating SmileyLlama should generally be thought of as converting Llama into a CLM, rather than simply adding an additional functionality to Llama.
S4 Supporting Tables
Table S1: Frequency of generating or sampling undesirable chemical structures in ChEMBL, SmileyLlama, and other SOTA methods. Here, we randomly sample 10,000 molecules from the ChEMBL dataset, SmileyLlama and our retrained models to compare their statistics of five different categories of undesirable chemical patterns.
Model ChEMBL SmileyLlama LSTM GPT S4
Unsaturated Benzene 0.55% 0.1 % 0.04% 0.04% 0.03%
Unsaturated Naphthalene 0.02% 0.01% 0.02% 0.01% 0.01%
Wrong Pyrrole 0.29% 0.02% 0.02% 0.03% 0.04%
Cyclopentadiene Ylidene 0.96% 0.1 % 0.04% 0.06% 0.05%
Benzyne $0$ 0.0% 0.0% 0.0%
Table S2: SMILES strings and scores for selected generated molecules. All Synthetic Accessibility (SA) Scores were calculated using the method by Ertl and Schuffenhauer ertlEstimationSyntheticAccessibility2009.
SMILES string Prompt Vina score SA Score
c1(-c2cc(-c3cc(cc(c3)F)F)nc(n2)N2C[C@H](O[C@H](C2)C)C)cc2c(nc(nc2N)N)cc1 SARS2Pro -9.6 3.44
c1(nc2cc(ccc2c(n1)Nc1cc2c(cc1F)nc(nc2N)N)F)Nc1cc2c(cc1F)[nH]nc2 SARS2Pro -9.9 2.95
c1(cc(-c2cc3c(nc(nc3Nc3cc4c(cc3F)OCCO4)N)cc2)cc2c1[nH]nc2)F SARS2Pro -10.4 2.82
c1(-c2cc3c(nc(nc3Nc3cc(ccc3)C(N3C[C@H](O[C@H](C)C3)C)=O)N)cc2)cc2c([nH]nc2)cc1 SARS2Pro -10.0 3.36
C1N(c2cc(-c3cc4c(nc(nc4N)N)cc3)nc(n2)Nc2cc3c(cc2F)occ3)CCOC1 SARS2Pro+Ro5 -9.6 2.86
c1cc2cc(-c3nc4c(ncnc4Nc4cc5c(cc4F)OCCO5)cc3)ccc2nc1N SARS2Pro+Ro5 -10.2 2.68
c1cc(-c2cc3c(nc(nc3Nc3cc4c(cc3F)OCCO4)N)cc2)cc2c1[nH]nc2C SARS2Pro+Ro5 -10.2 2.72
C1(N(c2cc(-c3nc4c(cccc4)c(n3)Nc3cc4c([nH]nc4)cc3)cc(c2)F)C(=O)CC1)=O SARS2Pro+Ro5 -10.1 2.64
Table S3: Performance of Meta-Llama-3.1-8B-Instruct versus SmileyLlama in a variety of LLM benchmarks. We report the mean score, assessed with the exact_match scoring method.
Model Llama-3.1-8B-Instruct SmileyLlama
MMLU (0-shot) 0.68 0.65
GPQA (chain of thought) 0.30 0.038
Math-Hard 0.24 0.039
MMLU-Pro 0.47 0.063
S5 Supporting Figures
Figure S1: Guacamol benchmark results as a function of temperature. We used n=3 runs to assess how validity, uniqueness and novelty depends on temperature, a parameter which affects the randomness of outputs generated during inference. Error bars are defined by the standard error. Temperature can be increased during inference for more random molecules, sacrificing validity for diversity and novelty. We generally use a temperature of T=1.0 during this study.
Figure S2: Validity, uniqueness and novelty over optimization for SARS2-Mpro inhibition We plot the validity (fraction of molecules which are valid), uniqueness (number of distinct molecules) and novelty (fraction of molecules which have not been generated before) of molecules sampled from the generative models of iMiner with 2000 molecules per epoch, and for SmileyLlama with DPO (SL+DPO) with 2000 and 1000 molecules per epoch. Note that novelty is not defined with respect to ChEMBL; here it’s a measure of how good the model is at generating molecules that it has not generated in a previous iteration.
Figure S3: SmileyLlama de novo generated molecules in the active site of SARS2 main protease. Surface rendering of the SmileyLlama generated molecules in the SARS2 Mpro canonical binding pocket. Generated by SmileyLlama after optimization with the SARS2PRO and SARS2Pro+Ro5 prompts. Table S3 provides their SMILES string and docking scores.
Figure S4: A visualization of how the SmileyLlama SFT procedure affects zero-shot Massive Multitask Language Understanding (MMLU) score75. Generally, MMLU scores appear to decrease with fine-tuning. One might expect that fine-tuning which improves the ability to generate molecules may translate to improved performance on chemistry, but this was found to be the opposite of the case; chemistry-specific sections of the MMLU saw among the worst degradation in score after SFT. Anecdotally, SmileyLlama will often respond to a chemistry-related query with a SMILES string. The sharpest decrease in performance was in the "Moral Scenarios" section. This may be due to much of the "moral knowledge" coming during post-training. The "moral knowledge" section of the MMLU had the largest difference in score between Between the base and instruct Llama models; while SmileyLlama scores about 41% compared to Llama-3.1-8B Instruct’s 57%, the base model Llama-3.1-8B score centers around 28%. On the other hand, there are some tasks, such as "Machine Learning", where performance appears to improve, although the standard error bars for each of these tasks include the mean so we cannot conclusively say that the SFT procedure used for SmileyLlama improved Llama’s abilities outside of SMILES string generation.

### v4

@@ 3.4 SmileyLlama Outside of Chemical Language Modeling
While SFT and DPO clearly alters Llama in the creation of SmileyLlama, one might wonder whether SmileyLlama still retains the full LLM knowledge of the Llama-3.1-8B model from which it derives. We find that SmileyLlama can still converse in English if it is prompted to do so, and some sample conversations are included in Section 3 in the Supplementary Information. However, we notice that its responses are often shorter than those of Llama-3.1-8B and can contain SMILES string non-sequiturs, which seem to be more likely when dealing with prompts relating to chemistry. As a more quantitative measure of its residual abilities, we evaluate performance of SmileyLlama on the Language Model Evaluation Harness containing the MMLU GPQA, Math-Hard, and MMLU-Pro benchmarkseval-harness, mmlu, wang2024mmlupro, rein2023gpqa, hendrycks2021math. Supplementary Table S3 and Figure S4 show that SmileyLlama generally performs worse on moral scenarios, and amusingly SmileyLlama also performs worse on subjects dealing with chemistry than Llama. This in part is likely due to the tendency for SmileyLlama to complete prompts relating to chemistry with a SMILES string as noted above. In addition, accuracy errors in the MMLU tests have also been noted recentlygema2025mmlu, and thus SmileyLlama’s degraded performance in chemistry may also be in part an artifact of poor testing benchmarks. Overall this result is somewhat encouraging, since it implies the possibility that LLM-derived CLMs can inherit and take advantage of the natural language processing ability of their foundation model. SmileyLlama already does this— we can steer the properties of the molecules it generates and the chemical space it explores using natural language prompts while still retaining some ability to process non-chemical natural language. However, more work will be required to view SmileyLlama as an additional capability of an LLM, as opposed to its current and primary benefit as a CLM.
@@
We have investigated the adaptation of LLMs to molecular chemistry problems that have been largely in the domain of CLMs. We find that the key advantage that LLMs have when performing as CLMs comes from prompting: information about the region of interest in chemical space can be specified in the prompt, letting the model generate chemical structures in this region. Furthermore, we showed that training an LLM on a variety of prompts allows for good performance in multiple CLM tasks, and found evidence to suggest that the LLM models can correctly generalize to scenarios with new prompts that leverage the abilities developed when trained on similar prompts. Concretely, we showed that Meta-Llama-3.1-8B-Instruct can be converted into a CLM on par with or exceeding the performance of modern CLMsmol-gen-bg3, mol-gen-bg4, mol-gen-gpt1, mol-opt1, iminer, rather than only demonstrating that LLMs can modify currently existing SMILES strings to have different properties.
@@
Our study clarifies a few crucial points for CLMs derived from LLMs going forward. First, it is not necessary to pretrain a specialized model on chemistry-specific text to generate molecules from a text description; a much less resource-intensive SFT training run on prompt-following on a dataset of a few million molecules with a commodity LLM is enough to do that. Second, DPO provides another resource-efficient way of optimizing the model to produce molecules that score well on a targeted objective without needing in-context examples, instead relying on the generative nature of the model itself for good and bad examples. Using these optimization methods, we showed that data specific to a target protein can generate increasingly better inhibitor molecules as determined by docking scores. A corollary to this is the finding that SmileyLlama can combine its knowledge gained during single-objective optimization to perform well at a task specifying multiple objectives, elicited by combining the prompts, which is a welcome outcome. SmileyLlama was able to generate molecules which satisfy Lipinski’s rule of 5 and a high docking score on SARS2-Mpro even though SmileyLlama was not trained on the generation of molecules with both prompts, but instead is found to generate molecules with both attributes at inference.
It should be mentioned that both high docking and drug-likeness scores are not enough for a molecule to be a good drug candidate; good candidates must also inhibit so-called "off-target effects", must be robust to mutation of the virus (generally through binding to essential residues), among other downstream requirements. While SmileyLlama is not being explicitly optimized for generating molecules with these qualities in this work, the DPO framework laid out here is extensible to optimizing molecules for these characteristics: for instance, off-target effects can be avoided by penalizing docking to proteins similar to the target. Furthermore, there are still limitations to and tradeoffs within the SmileyLlama framework and our investigation. While DPO improves adherence to the prompt, it does so at the cost of narrowing the distribution of properties or diversity, which may not be desirable in all application areas outside of discovery. Furthermore, SmileyLlama still struggles in data-poor regimes, for instance in the task of generating macrocycles.
@@
The prompting and optimization framework for modifying LLMs to explore chemical space broadly or to narrow the search to specific regions shown here could also be leveraged for molecular design outside of drug discovery, such as the use of SMILES for elaborating on transition metal complexesRasmussen2025. One could also imagine that casting a chemical problem as a linguistic construct could also allow for new applications, such as our recent work on chemical synthesisSun2025. As with many of the fields touched by LLMs this decade, the newly opened frontier of possibility in chemistry is as vast as it is exciting.
@@ Acknowledgments
Code used for this study can be found at https://anonymous.4open.science/r/SmileyLlama-gh-D038/. SmileyLlama and its derivatives optimized with DPO can be found at https://huggingface.co/collections/ THGLab/smileyllama-6880b14e4e4d708001564062